OpenStudy (anonymous):

A heated piece of metal cools according to the function c(x) = (.5)x − 7, where x is measured in hours. A device is added that aids in cooling according to the function h(x) = −x − 2. What will be the temperature of the metal after two hours? −4° Celsius 28° Celsius 32° Celsius 38° Celsius

OpenStudy (anonymous):

@texaschic101 @FriedRice

OpenStudy (anonymous):

@DominicNg @IMStuck

OpenStudy (anonymous):

hmmm c(x) and h(x) are given in the problem statement. Just copy them into C(x)=c(x)+h(x).

OpenStudy (anonymous):

Then, C(x) = (.5)^x - x - 9 = (.5)^x - x - 9 Now, evaluate this function at x = 2.

OpenStudy (anonymous):

If C(x) = (.5)^2 - x - 9, then C(2) = (.5)^2 - 11 = (1/4)-11.

OpenStudy (anonymous):

are you unclear of whats happening?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait brb

OpenStudy (anonymous):

ok remember i said you can just copy the two equations into C(x)=c(x)+h(x).

OpenStudy (anonymous):

do that